OpenStudy (anonymous):

I need help with finding the x and y how would you solve; 2x+3y=4 4x+6y=8

OpenStudy (anonymous):

by elimination u subtract the second term from the first term ( remember that when u subtract all the signs change to the opposite)

OpenStudy (anonymous):

Multiply the first equation by -2. -2(2x+3y=4) 4x+6y=8 -4x - 6y = -8 4x + 6y = 8 0 = 0 THis is an identity outcome. Identity outcomes mean infinitely many solutions.